Two long, parallel wires are attracted to each other by a force per unit length of 440 𝜇𝑁/𝑚


when they are separated by a vertical distance of 0.80 m. The current in the upper wire is


13.0 A to the right. Determine the location of the line in the plane of the two wires along


which the total magnetic field is zero.



1
Expert's answer
2022-03-30T13:50:56-0400

Explanation


"\qquad\qquad
\begin{aligned}
\small \frac{F}{\Delta L}&=\small \mu.\frac{i_1i_2}{2\pi d} = F_{given}\\
\small i_2&=\small F_g.\frac{(2\pi d)}{\mu.i_1}\\
\\
\small B_{total}&=\small \mu.\frac{i_1}{2\pi x}+\mu.\frac{i_2}{2\pi(d-x)}= 0\\
\small \frac{i_1}{x}&=\small \frac{i_2}{x-d}\\
\small x&=\small \frac{i_1 d}{(i_1-i_2)}
\end{aligned}"


  • Here "\small i_1" is the 13 A current on the top line and the current on the other line can be calculated.
  • "\small d" is the separation between the wires.
  • Then the distance to the layer where the magnetic field is zero can be calculated.
  • "\small i_2>i_1" should be as the wires are held by an attracting force meaning that both currents are along the same direction.
  • Then probably the x should receive a negative value meaning that it is not in between the wires but outside of them.